### v1

# Character tables are ideal Perron similarities

Abstract
An invertible matrix is called a Perron similarity if it diagonalizes an irreducible, nonnegative matrix. Each Perron similarity gives a nontrivial polyhedral cone, called the spectracone, and polytope, called the spectratope, of realizable spectra (thought of as vectors in complex Euclidean space). A Perron similarity is called ideal if its spectratope coincides with the conical hulls of its rows. Identifying ideal Perron similarities is of great interest in the pursuit of the longstanding nonnegative inverse eigenvalue problem.
In this work, it is shown that the character table of a finite group is an ideal Perron similarity. In addition to expanding ideal Perron similarities to include a broad class of matrices, the results unify previous works into a single, theoretical framework.
It is demonstrated that the spectracone can be described by finitely-many group-theoretic inequalities. When the character table is real, we derive a group-theoretic formula for the volume of the projected Perron spectratope, which is a simplex. Finally, an implication for further research is given.
keywords:
character , character table , group , nonnegative matrix , Perron similarity , polyhedral cone , polytope , representation
MSC:
[2020] 15A29 , 20C15 , 20C99 , 15B51 , 52B99
[…]
[add]organization=Division of Engineering & Mathematics, University of Washington Bothell,addressline=18115 Campus Way NE, city=Bothell, postcode=98011-8246, state=WA, country=U.S.A
1 Introduction
The nonnegative inverse eigenvalue problem (NIEP) is to determine which multisets (hereinafter, lists) of complex numbers occur as the spectra of entry-wise nonnegative matrices. The NIEP has proven to be one of the most challenging and sought-after problems in matrix analysis and includes a number of subproblems that are equally as daunting [8].
If $A$ is an $n$-by-$n$ entrywise nonnegative matrix with spectrum $\Lambda=\{\lambda_{1},\ldots,\lambda_{n}\}$, then the list $\Lambda$ is called realizable, and $A$ is called a realizing matrix (for $\Lambda$). In such a case,
$\displaystyle\rho\left(\Lambda\right)$ $\displaystyle\coloneqq\max_{1\leqslant k\leqslant n}\left\{|\lambda_{k}|\right\}\in\Lambda,$ (1.1)
$\displaystyle\overline{\Lambda}$ $\displaystyle\coloneqq\left\{\overline{\lambda_{1}},\dots,\overline{\lambda_{n}}\right\}=\Lambda,$ (1.2)
[…]
and
$\displaystyle\left[s_{k}(\Lambda)\right]^{\ell}$ $\displaystyle\leqslant n^{\ell-1}s_{k\ell}(\Lambda),\forall k,\ell\in\mathbb{N}$
are the well-known necessary conditions.
If $S\in\mathsf{GL}_{n}(\mathbb{F})$, where $\mathbb{F}\in\{\mathbb{R},\mathbb{C}\}$, and $e$ denotes the all-ones vector of appropriate size, then the polyhedral cone
$\mathcal{C}(S):=\{x\in\mathbb{F}^{n}\mid M_{x}\coloneqq SD_{x}S^{-1}\geqslant 0\}\supseteq\{\alpha e\mid\alpha\geqslant 0\},$
is called the (Perron) spectracone of $S$, and the polytope
$\mathcal{P}(S):=\left\{x\in\mathcal{C}(S)\mid M_{x}e=e\right\}\supseteq\{e\},$
is called the (Perron) spectratope of $S$. If there is a vector $x\in\mathbb{F}^{n}$ such that the matrix $M_{x}$ is irreducible and nonnegative, then $S$ is called a Perron similarity. The matrix $S$ is called ideal if $\mathcal{C}(S)=\mathcal{C}_{r}(S)$, where $\mathcal{C}_{r}(S)$ denotes the conical hull of the rows of $S$.
Identifying ideal Perron similarities is and has been of interest in the study of the NIEP [9, 10, 11]. In particular:
  • A matrix $H=[h_{ij}]\in\mathsf{M}_{n}(\mathbb{R})$ is called a Hadamard matrix if $h_{ij}\in\{\pm 1\}$ and $HH^{\top}=nI_{n}$. A Hadamard matrix $H$ is called normalized if $He_{1}=e$ and $e_{1}^{\top}H=e^{\top}$. Johnson and Paparella [9] used the theory of association schemes to show that the Walsh matrix
    $H_{2^{n}}\coloneqq\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}^{\otimes n},\ n\in\mathbb{N}_{0}$ (1.4)
    is ideal. Although every normalized Hadamard matrix is a Perron similarity, not every such matrix is ideal [11, Remark 4.31].
  • A matrix $H=[h_{ij}]\in\mathsf{M}_{n}(\mathbb{C})$ is called a complex Hadamard matrix if $|h_{ij}|=1$ and $HH^{\ast}=nI_{n}$. A complex Hadamard matrix $H$ is called dephased if $He_{1}=e$ and $e_{1}^{\top}H=e^{\top}$.
    In a more recent work, Johnson and Paparella [11, Corollary 7.1] showed that the non-normalized discrete Fourier transform matrix
    $F=F_{n}=[\omega_{n}^{(i-1)(j-1)}],\ \omega_{n}\coloneqq e^{\frac{2\pi}{n}\mathsf{i}},$ (1.5)
    a dephased complex Hadamard matrix, is also ideal by appealing to the fact that $F_{n}$ is a Vandermonde matrix with respect to the zeros of the polynomial $t^{n}-1$. Furthermore, any arbitrary Kronecker product of the form
    $\left(\bigotimes_{j=1}^{N}F_{n_{j}}\right)\otimes H_{2^{k}},\ k\in\mathbb{N}_{0},\ n_{j}\in\mathbb{N},\ 1\leqslant j\leqslant N,$ (1.6)
    is also ideal [11, Corollary 7.17].
If $H$ is a dephased complex Hadamard matrix, then $H$ is a Perron similarity because
$\frac{1}{n}H(e_{1}e_{1}^{\top})H^{\ast}=\frac{1}{n}ee^{\top}>0,$
but a dephased complex Hadamard matrix need not be ideal: notice that there are values $\theta\in[0,\pi)$ such that the multisets
$\{1,\mathsf{i}e^{\mathsf{i}\theta},-1,-\mathsf{i}e^{\mathsf{i}\theta}\}$
and
$\{1,-\mathsf{i}e^{\mathsf{i}\theta},-1,\mathsf{i}e^{\mathsf{i}\theta}\}$
fail the self-conjugacy condition (1.2). Consequently, there are values $\theta\in[0,\pi)$ such that the dephased complex Hadamard matrix
$F_{4}^{(1)}(\theta)\coloneqq\begin{bmatrix}1&1&1&1\\ 1&\mathsf{i}e^{\mathsf{i}\theta}&-1&-\mathsf{i}e^{\mathsf{i}\theta}\\ 1&-1&1&-1\\ 1&-\mathsf{i}e^{\mathsf{i}\theta}&-1&\mathsf{i}e^{\mathsf{i}\theta}\end{bmatrix},\ \theta\in[0,\pi).$
is not ideal.
It is thus natural to ask if the matrices given by equations (1.4)–(1.6) belong to a larger class of ideal Perron similarities. Interestingly, these matrices are character tables, which leads naturally to the question: is every character table an ideal Perron similarity?
We utilize character theory to affirmatively answer the question posed above. This not only expands ideal Perron similarities to a broad class of matrices but also unifies results that appeared previously in the literature [3, 9, 10, 11] within a single theoretical framework.
Constructive results are highly sought-after in the NIEP. Indeed, according to Chu [2]: Very few of these theoretical results are ready for implementation to actually compute [the realizing] matrix. The most constructive result we have seen is the sufficient condition studied by Soules [175]. But the condition there is still limited because the construction depends on the specification of the Perron vector—in particular, the components of the Perron eigenvector need to satisfy certain inequalities in order for the construction to work. [2, p. 18]. Unfortunately, what Chu wrote in the late 1990s still rings true—for example, a constructive version of a result due to Fiedler that every Suleĭmanova spectrum is symmetrically realizable is only known for low dimensions and Hadamard orders [9].
The results contained here are constructive: every character table yields a polytope of realizable spectra that is closed with respect to the Hadamard product. Of particular interest are the extreme points of the polytope corresponding to the character table of a finite Abelian group as they are located on the boundary of the set of realizable spectra.
2 Background & Notation
For ease of notation, $\mathbb{N}$ denotes the set of positive integers and $\mathbb{N}_{0}\coloneqq\mathbb{N}\cup\{0\}$. If $n\in\mathbb{N}$, then $\left[n\right]\coloneqq\{k\in\mathbb{N}\mid 1\leqslant k\leqslant n\}$.
The set of $m$-by-$n$ matrices with entries from a field $\mathbb{F}$ is denoted by $\mathsf{M}_{m\times n}(\mathbb{F})$. If $m=n$, then $\mathsf{M}_{n\times n}(\mathbb{F})$ is abbreviated to $\mathsf{M}_{n}(\mathbb{F})$. The set of nonsingular matrices in $\mathsf{M}_{n}(\mathbb{F})$ is denoted by $\mathsf{GL}_{n}\left(\mathbb{F}\right)$.
If $x\in\mathbb{F}^{n}$, then $x_{k}$ or $[x]_{k}$ denotes the $k\textsuperscript{th}$-entry of $x$ and $D_{x}\in\mathsf{M}_{n}$ denotes the diagonal matrix whose $(k,k)$-entry is $x_{k}$. A vector $x\in\mathbb{F}^{n}$ is called totally nonzero if $x_{k}\neq 0,\ \forall k\in\left[n\right]$. If $x\in\mathbb{F}^{n}$ is totally nonzero, then $(D_{x})^{-1}=D_{x^{-1}}$, where $x^{-1}$ denotes the entrywise inverse of $x$ (i.e., the inverse with respect to the Hadamard product).
By the mechanics of matrix multiplication, if $S\in\mathsf{GL}_{n}(\mathbb{F})$ and $t_{ij}$ denotes the $(i,j)$-entry of $S^{-1}$, then
$\left[SD_{x}S^{-1}\right]_{ij}=\sum_{k=1}^{n}s_{ik}(x_{k}t_{kj})=\sum_{k=1}^{n}(s_{ik}t_{kj})x_{k}.$ (2.1)
Denote by $I$, $e$, $e_{k}$, and $0$ the identity matrix, the all-ones vector, the $k\textsuperscript{th}$ canonical basis vector, and the zero vector, respectively. The size of each aforementioned object is implied by its context.
If $A\in\mathsf{M}_{m\times n}(\mathbb{F})$, then:
  • $a_{ij}$, $a_{i,j}$, or $[A]_{ij}$ denotes the $(i,j)$-entry of $A$;
  • $A^{\top}$ denotes the transpose of $A$;
  • $\overline{A}=[\overline{a_{ij}}]$ denotes the (entrywise) conjugate of $A$;
  • $A^{\ast}\coloneqq\overline{A^{\top}}=\overline{A}^{\top}$ denotes the conjugate-transpose of $A$;
  • $r_{i}(A)\coloneqq A^{\top}e_{i}$ denotes the $i\textsuperscript{th}$-row of $A$ as a column vector and when the context is clear, $r_{i}(A)$ is abbreviated to $r_{i}$; and
  • when $m=n$, $\operatorname{\mathsf{spec}}A=\operatorname{\mathsf{spec}}(A)$ denotes the spectrum of $A$ and $\rho=\rho(A)$ denotes the spectral radius of $A$.
If $X\subseteq\mathbb{F}^{n}$, then the conical hull of $X$, denoted by $\operatorname{\mathsf{coni}}X=\operatorname{\mathsf{coni}}(X)$, is defined by
$\operatorname{\mathsf{coni}}{X}=\begin{cases}\{0\},&X=\varnothing\\ \left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,~\alpha_{k}\geqslant 0\right\},&X\neq\varnothing.\end{cases}$
i.e., when $X$ is nonempty, $\operatorname{\mathsf{coni}}X$ consists of all conical combinations. Similarly, the convex hull of $X$, denoted by $\operatorname{\mathsf{conv}}{X}=\operatorname{\mathsf{conv}}(X)$ is defined by
$\operatorname{\mathsf{conv}}{X}=\begin{cases}\{0\},&X=\varnothing\\ \left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,\ \sum_{k=1}^{m}\alpha_{k}=1,\ \alpha_{k}\geqslant 0\right\},&X\neq\varnothing\end{cases}.$
The conical hull or convex hull of a finite list $\{x_{1},\ldots,x_{n}\}$ is abbreviated to $\operatorname{\mathsf{coni}}(x_{1},\ldots,x_{n})$ or $\operatorname{\mathsf{conv}}(x_{1},\ldots,x_{n})$, respectively.
2.1 Characters & Representations
Hereinafter, $G=(G,e_{G},\cdot)$ is a finite group. For ease of notation, $a\cdot b$ is abbreviated to $ab$. If $g\in G$, then $\operatorname{\mathsf{cl}}(g)\coloneqq\{aga^{-1}\mid a\in G\}$ denotes the conjugacy class of $g$ and $C_{G}(g)\coloneqq\{a\in G\mid ag=ga\}$ denotes the centralizer of $g$ (in $G$). Because $\{\operatorname{\mathsf{cl}}(g)\mid g\in G\}$ forms a partition of $G$, we write $x\sim y$ whenever $x,y\in\operatorname{\mathsf{cl}}(g)$.
If $\rho:G\longrightarrow\textsf{GL}_{n}(\mathbb{C})$ is a homomorphism, then $\rho$ is called a (matrix) representation (of $G$). As $\rho$ is a homomorphism, $\rho(e_{G})=I_{n}$ and $\rho(g^{-1})=\rho(g)^{-1}$.
If $\rho$ and $\sigma$ are representations, then:
  • the conjugate transpose of $\rho$, denoted by $\rho^{\ast}$, is the representation defined by $\rho^{\ast}(g)=\rho(g)^{\ast},\ \forall g\in G$;
  • the direct sum of $\rho$ and $\sigma$, denoted by $\rho\oplus\sigma$, is the representation defined by $(\rho\oplus\sigma)(g)=\rho(g)\oplus\sigma(g),\ \forall g\in G$;
  • the Kronecker or tensor product of $\rho$ and $\sigma$, denoted by $\rho\otimes\sigma$, is the representation defined by $(\rho\otimes\sigma)(g)=\rho(g)\otimes\sigma(g),\ \forall g\in G$; and
  • $\rho$ and $\sigma$ are isomorphic or similar if there is an invertible matrix $S$ such that $\rho(g)=S\sigma(g)S^{-1},\ \forall g\in G$.
For a representation $\rho$, the function $\chi_{\rho}:G\longrightarrow\mathbb{C}$, defined by
$\chi_{\rho}(g)=\text{tr}(\rho(g)),\ \forall g\in G,$
[…]
The following properties are well known (see, e.g., Fulton and Harris [5, p. 13] or Serre [15, Propositions 1 and 2]):
$\displaystyle\chi_{\rho\oplus\sigma}(g)$ $\displaystyle=\chi_{\rho}(g)+\chi_{\sigma}(g)$
$\displaystyle\chi_{\rho\otimes\sigma}(g)$ $\displaystyle=\chi_{\rho}(g)\cdot\chi_{\sigma}(g)$ (2.2)
$\displaystyle\chi_{\rho^{\ast}}(g)$ $\displaystyle=\overline{\chi_{\rho}(g)}=\chi_{\rho}(g^{-1})$ (2.3)
$\displaystyle\chi_{\rho}(g)$ $\displaystyle=\chi_{\rho}(h),\ g\sim h$ (2.4)
Equation 2.4 ensures that $\chi_{\rho}$ is a class function.
A subspace $W$ of $\mathbb{C}^{n}$ is called $G$-invariant or a $G$-invariant subspace if
$\rho(g)w\in W,\ \forall g\in G,\ \forall w\in W.$
Obviously, the trivial subspaces $\{0\}$ and $\mathbb{C}^{n}$ are $G$-invariant; if $\mathbb{C}^{n}$ has a proper $G$-invariant subspace, then $\rho$ is called reducible or a reducible representation. If $\mathbb{C}^{n}$ has no proper $G$-invariant subspace, then $\rho$ is called irreducible or an irreducible representation. A character of an irreducible representation is called an irreducible character.
If $\rho$ is a reducible representation, then there is an invertible matrix $S$ such that
$S^{-1}\rho(g)S=\begin{bmatrix}\rho_{1}(g)&\sigma(g)\\ 0&\rho_{2}(g)\end{bmatrix}$,
where $\rho_{1}=\left.\rho\right|_{W}$, $\rho_{2}$ is the representation of $G$ on $\mathbb{C}^{n}/W$, and $\sigma(g)$ is a rectangular matrix [4, p. 848].
The representation $g\in G\longmapsto I_{n}\in\mathsf{GL_{n}(\mathbb{C})}$ is called the ($n$-dimensional) trivial representation and is irreducible when and only when $n=1$.
If $\rho$ and $\sigma$ are representations, then
$\langle\chi_{\rho},\chi_{\sigma}\rangle\coloneqq\frac{1}{|G|}\sum_{g\in G}\overline{\chi_{\rho}(g)}\chi_{\sigma}(g)=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\left(\overline{\chi_{\rho}(g_{k})}\chi_{\sigma}(g_{k})\right)$
defines an inner product on characters (see, e.g., Artin [1, pp. 300]).
The number of distinct irreducible representations (up to isomorphism) of a finite group equals the number of its distinct conjugacy classes (see, e.g., Artin [1, Theorem 10.4.6(b)]; Fulton and Harris [5, Corollary 2.13]; or Serre [15, Theorem 7]). The following well-known theorem (see, e.g., Artin [1, Theorem 10.4.6(a) and Corollary 10.4.8] or Dummit and Foote [4, Theorem 15]) will be crucial in what follows.
Theorem 2.1.
[…]
charlinearcombo If $\rho_{1},\ldots,\rho_{n}$ are the distinct irreducible representations of $G$ and $\chi_{1},\ldots,\chi_{n}$ are their respective characters, then
$\langle\chi_{i},\chi_{j}\rangle=\delta_{ij}.$
If $\rho$ is a representation, then
$\chi_{\rho}(g)=\sum_{k=1}^{n}\langle\chi_{\rho},\chi_{k}\rangle\chi_{k}(g),\ \forall g\in G.$ (2.5)
Every character is, in fact, a nonnegative integral linear combination of the irreducible characters [4, p. 868].
2.2 Character Tables
Suppose that $\rho_{1},\ldots,\rho_{n}$ are, up to isomorphism, the distinct irreducible representations of $G$ and $\operatorname{\mathsf{cl}}(g_{1}),\ldots,\operatorname{\mathsf{cl}}(g_{n})$ are the distinct conjugacy classes of $G$. For ease of notation, $\chi_{\rho_{k}}$ is abbreviated to $\chi_{k}$. The $n$-by-$n$ matrix $Q$
$Q=\begin{bmatrix}\chi_{1}(g_{1})&\cdots&\chi_{1}(g_{n})\\ \vdots&\ddots&\vdots\\ \chi_{n}(g_{1})&\cdots&\chi_{n}(g_{n})\end{bmatrix}$,
i.e., $q_{ij}\coloneqq\chi_{i}(g_{j})$, is called the character table of $G$. The class function property (2.4) ensures that the matrix $Q$ is unique up to permutation similarity.
Let $\rho_{1}:G_{1}\longrightarrow\mathsf{GL}_{n_{1}}(\mathbb{C})$ and $\rho_{2}:G_{2}\longrightarrow\mathsf{GL}_{n_{2}}(\mathbb{C})$ be representations of finite groups $G_{1}$ and $G_{2}$, respectively. The Kronecker product of matrices defines a representation of $G_{1}\times G_{2}$, also called the tensor product, via
$(\rho_{1}\otimes\rho_{2})(g_{1},g_{2})=\rho_{1}(g_{1})\otimes\rho_{2}(g_{2}),\ \forall(g_{1},g_{2})\in G_{1}\times G_{2}.$
If, in addition, $\rho_{1}$ and $\rho_{2}$ are irreducible, then $\rho_{1}\otimes\rho_{2}$ is an irreducible representation of $G_{1}\times G_{2}$ [15, Theorem 10(i)]. Conversely, every irreducible representation of $G_{1}\times G_{2}$ is isomorphic to a tensor product $\rho_{1}\otimes\rho_{2}$, where $\rho_{1}$ and $\rho_{2}$ are irreducible representations of $G_{1}$ and $G_{2}$, respectively [15, Theorem 10(ii)]. Hence, if $Q_{1}$ and $Q_{2}$ are character tables of $G_{1}$ and $G_{2}$, respectively, then $Q_{1}\otimes Q_{2}$ is the character table of $G_{1}\times G_{2}$. This result generalizes as follows: if $Q_{1},\ldots,Q_{m}$ are character tables of $G_{1},\ldots,G_{m}$, respectively, then $\bigotimes_{k=1}^{m}Q_{k}$ is the character table of $\prod_{k=1}^{m}G_{k}$.
2.3 Nonnegative & Stochastic Matrices
If $A\in\mathsf{M}_{n}(\mathbb{C})$, $n\geqslant 2$, then $A$ is called reducible if there is a permutation matrix $P$ such that
$\displaystyle P^{\top}AP=\begin{bmatrix}A_{11}&A_{12}\\ 0&A_{22}\end{bmatrix}$,
where $A_{11}$ and $A_{22}$ are non-empty square matrices. If $A$ is not reducible, then A is called irreducible. Clearly, entrywise positive matrices are irreducible.
If $x\in\mathbb{R}^{n}$ and $x_{i}\geqslant 0,\ \forall i\in\left[n\right]$ ($x_{i}>0,\ \forall i\in\left[n\right]$), then is called nonnegative (respectively, positive), and we write $x\geqslant 0$ (respectively, $x>0$). A similar definition applies to real matrices. If $x\in\mathbb{R}^{n}$ ($A\in\mathsf{M}_{n}(\mathbb{R})$), then $x$ (respectively, $A$) is viewed as a complex vector (respectively, matrix) via the map $x\longmapsto x+0\mathsf{i}$ (respectively, $A\longmapsto A+0\mathsf{i}$).
If $A\geqslant 0$ and
$\sum_{j=1}^{n}a_{ij}=1,~\forall~i\in\left[n\right],$
then $A$ is called (row) stochastic. If $A\geqslant 0$, then $A$ is stochastic if and only if $Ae=e$. Furthermore, if $A$ is stochastic, then $1\in\operatorname{\mathsf{spec}}(A)$ and $\rho(A)=1$. It is known that the NIEP and the stochastic NIEP are equivalent (see, e.g., Johnson [7, p. 114]).
3 Spectral Polyhedra & Perron Similarities
Definition 3.1.
[…]
spectratope If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then:
  1. (i)
    $\mathcal{C}(S)\coloneqq\{x\in\mathbb{C}^{n}\mid M_{x}=M_{x}(S)\coloneqq SD_{x}S^{-1}\geqslant 0\}$ is called the (Perron) spectracone of $S$;
  2. (ii)
    $\mathcal{P}(S)\coloneqq\{x\in\mathcal{C}(S)\mid M_{x}e=e\}$ is called the (Perron) spectratope of $S$; and
  3. (iii)
    $\mathcal{A}(S)\coloneqq\{M_{x}\in\mathsf{M}_{n}(\mathbb{R})\mid x\in\mathcal{C}(S)\}$.
Since $M_{e}=SD_{e}S^{-1}=SIS^{-1}=I\geqslant 0$ and $I$ is stochastic, it follows that $e\in\mathcal{P}(S)\subset\mathcal{C}(S)$ and all three sets are nonempty. If $\mathcal{C}(S)=\operatorname{\mathsf{coni}}(e)$, $\mathcal{P}(S)=\{e\}$, or $\mathcal{A}(S)=\operatorname{\mathsf{coni}}(I_{n})$, then $\mathcal{C}(S)$, $\mathcal{P}(S)$, and $\mathcal{A}(S)$ are called trivial; otherwise, they are called nontrivial.
It is known that $\mathcal{C}(S)$ is a polyhedral cone and $\mathcal{P}(S)$ is a polytope and both sets are closed with respect to the Hadamard product [11, Theorems 4.4 and 4.6]. The set $\mathcal{A}(S)$ is a cone that is closed with respect to matrix multiplication [11, Theorem 4.4(iii)].
Reall from the introduction that an invertible matrix $S$ is called ideal if $\mathcal{C}(S)=\mathcal{C}_{r}(S)$, where $\mathcal{C}_{r}(S)$ denotes the conical hull of the rows of $S$. It is known that $S$ is ideal if and only if $e\in\mathcal{C}_{r}(S)$ and $r_{i}\in\mathcal{C}(S),\ \forall i\in\left[n\right]$, i.e., every row, viewed as a column vector, belongs to its spectracone [11, Theorem 4.21].
Building on a definition that previously appeared for real matrices [10, Definition 2.3], we now extend the concept of row Hadamard conic to complex matrices.
Definition 3.2.
[…]
rhc If $S\in\mathsf{M}_{m\times n}({\mathbb{C}})$, then $S$ is called row Hadamard conic (RHC) if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall i,j\in\left[m\right]$.
Proposition 3.3 ([11, Proposition 4.16]).
[…]
simple If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $x^{\top}S^{-1}\geqslant 0$ if and only if $x\in\mathcal{C}_{r}(S)$.
Proposition 3.4.
[…]
rhcrows If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $r_{i}\in\mathcal{C}(S)$ if and only if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),\ \forall j\in\left[n\right]$.
Proof.
Because the $j$th-row of $SD_{r_{i}}$ is $r_{j}\circ r_{i}$, it follows that
$\displaystyle r_{i}\in\mathcal{C}(S)$ $\displaystyle\Longleftrightarrow SD_{r_{i}}S^{-1}\geqslant 0$
$\displaystyle\Longleftrightarrow(SD_{r_{i}})S^{-1}\geqslant 0$
$\displaystyle\Longleftrightarrow\left(r_{j}\circ r_{i}\right)S^{-1}\geqslant 0,~\forall j\in\left[n\right]$
$\displaystyle\Longleftrightarrow r_{j}\circ r_{i}=r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall j\in\left[n$
as desired. ∎
[…]
Theorem 3.5.
[…]
RHC ideal If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $S$ is ideal if and only if $e\in\mathcal{C}_{r}(S)$ and $S$ is RHC.
If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $S$ is called a Perron similarity if there is a diagonal matrix $D$ such that $A=SDS^{-1}$ is irreducible and non-negative. If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $S$ is a Perron similarity if and only if there is a unique positive integer $k\in\left[n\right]$ such that $Se_{k}=\alpha x$ and $e_{k}^{\top}S^{-1}=\beta y^{\top}$, where $\alpha$ and $\beta$ are complex numbers such that $\alpha\beta>0$, and $x$ and $y$ are positive vectors [11, Theorem 4.14].
Lemma 3.6.
[…]
epairconstruct Supoose that $S\in\mathsf{GL}_{n}(\mathbb{C})$, $\lambda\in\mathbb{C}^{n}$, and $A\coloneqq SD_{\lambda}S^{-1}$. If $\exists k\in[n]$ such that $v\coloneqq Se_{k}$ is totally nonzero, then $(\lambda_{k},e)$ is an eigenpair of $D_{v^{-1}}AD_{v}$.
Proof.
Notice that
$\displaystyle(D_{v^{-1}}AD_{v})e$ $\displaystyle=D_{v^{-1}}((SD_{\lambda}S^{-1})v)$
$\displaystyle=D_{v^{-1}}((SD_{\lambda}S^{-1})Se_{k})$
$\displaystyle=D_{v^{-1}}((SD_{\lambda})e_{k})$
$\displaystyle=D_{v^{-1}}(\lambda_{k}v)$
$\displaystyle=\lambda_{k}e,$
and the result is established. ∎
If $S\in\mathsf{GL}_{n}(\mathbb{C})$ and $v>0$, then $\mathcal{C}(S)=\mathcal{C}(D_{v}S)$ [11, Theorem 4.10(iii)]. However, matrix transformations with respect to spectratopes remain unexplored. The following result is novel.
Theorem 3.7.
[…]
P(S)contain If $S\in\mathsf{GL}_{n}(\mathbb{C})$ is a Perron similarity, then there exists $k\in[n]$ such that $\mathcal{P}(S)\subseteq\mathcal{P}(D_{v^{-1}}S)$, where $v\coloneqq Se_{k}$.
Proof.
As $S$ is a Perron similarity, there is a positive integer $k\in[n]$, a nonzero complex number $\alpha$, and a positive vector $x$ such that $v\coloneqq Se_{k}=\alpha x$ is totally nonzero.
[…]
Example 3.8.
[…]
$SD_{x}S^{-1}=\begin{bmatrix}\frac{x_{1}+x_{2}}{2}&\frac{x_{1}-x_{2}}{4}\\ x_{1}-x_{2}&\frac{x_{1}+x_{2}}{2}\end{bmatrix}$
is stochastic if and only if $x_{1}=x_{2}=1$, i.e., $\mathcal{P}(S)$ is trivial. However, the matrix $H_{2}=\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}$ is ideal.
4 Main Results
In what follows, $Q$ is the character table of a finite group $G$. Since character tables are unique up to permutation similarity, it is assumed that $g_{1}=e_{G}$ and $\rho_{1}$ is the one-dimensional trivial representation.
Theorem 4.1.
The matrix $Q$ is a Perron similarity.
Proof.
Since
$\sum_{k=1}^{n}\overline{\chi_{k}(g)}\chi_{k}(h)=\begin{cases}|C_{G}(g)|,&g\sim h\\ 0,&g\nsim h\end{cases}$
(see, e.g., Dummit and Foote [4, Theorem 16]), it follows that
$\langle Qe_{j},Qe_{i}\rangle=(Qe_{i})^{\ast}Qe_{j}=\sum_{k=1}^{n}\overline{\chi_{k}(g_{i})}\chi_{k}(g_{j})=\begin{cases}|C_{G}(g_{i})|,&i=j\\ 0,&i\neq j\end{cases}$
i.e., the columns of $Q$ are orthogonal. Moreover, the matrix $D\coloneqq Q^{*}Q$ is an invertible diagonal matrix such that
$d_{ij}=\begin{cases}|C_{G}(g_{i})|,&i=j\\ 0,&i\neq j.\end{cases}$
Thus, $Q^{-1}=D^{-1}Q^{\ast}$ and
$\left[Q^{-1}\right]_{ij}=\frac{\overline{q_{ji}}}{|C_{G}(g_{i})|}$ (4.1)
Since $g_{1}=e_{G}$, we have
[…]
Via Equation 2.1,
$\displaystyle\left[M_{e_{1}}(Q)\right]_{ij}$ $\displaystyle=\left[QD_{e_{1}}Q^{-1}\right]_{ij}$
$\displaystyle=\sum_{k=1}^{n}\left(q_{ik}\left[Q^{-1}\right]_{kj}\right)[e_{1}]_{k}$
$\displaystyle=\frac{q_{i,1}\overline{q_{j,1}}}{|C_{G}(g_{1})|}$
$\displaystyle=\frac{\dim(\rho_{i})\dim(\rho_{j})}{|C_{G}(g_{1})|}>0.$
Thus, $M_{e_{1}}(Q)>0$. ∎
Theorem 4.2.
[…]
extremerayQ The matrix $Q$ is ideal.
Proof.
Since
[…]
it follows that $e_{1}^{\top}Q=e^{\top}$, i.e., $e\in\mathcal{C}_{r}(Q)$.
Suppose that $i,j\in[n]$. For ease of notation, let
$\alpha_{k}\coloneqq\langle\chi_{\rho_{i}\otimes\rho_{j}},\chi_{k}\rangle,\ \forall k\in\left[n\right]$
and $\alpha\coloneqq\begin{bmatrix}\alpha_{1}&\cdots&\alpha_{n}\end{bmatrix}^{\top}$. With the above in mind,
$\displaystyle r_{i}^{\top}\circ r_{j}^{\top}$ $\displaystyle=\begin{bmatrix}\chi_{i}(g_{1})\chi_{j}(g_{1})&\cdots&\chi_{i}(g_{n})\chi_{j}(g_{n})\end{bmatrix}$
$\displaystyle=\begin{bmatrix}\chi_{\rho_{i}\otimes\rho_{j}}(g_{1})&\cdots&\chi_{\rho_{i}\otimes\rho_{j}}(g_{n})\end{bmatrix}$ [Eq. 2.2]
$\displaystyle=\begin{bmatrix}\sum_{k=1}^{n}\alpha_{k}\chi_{k}(g_{1})&\cdots&\sum_{k=1}^{n}\alpha_{k}\chi_{k}(g_{n})\end{bmatrix}$ [Eq. 2.5]
$\displaystyle=\alpha^{\top}Q.$
[…]
Theorem 4.3.
[…]
ineqredux If $x\in\mathbb{C}^{n}$, then $M_{x}(Q)\geqslant 0$ if and only if
$\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})x_{k}\geqslant 0,\hskip 10.0pt\forall i\in[n].$ (4.2)
Proof.
First, recall that
$|G|=|C_{G}(g)||\operatorname{\mathsf{cl}}(g)|,\ \forall g\in G$ (4.3)
(see, e.g., Artin [1, p. 196]). By equations 2.1, 4.1, and 4.3,
$\displaystyle\left[M_{x}(Q)\right]_{ij}$ $\displaystyle=\sum_{k=1}^{n}\left(q_{ik}\left[Q^{-1}\right]_{kj}\right)x_{k}$
$\displaystyle=\sum_{k=1}^{n}\frac{\chi_{i}(g_{k})\overline{\chi_{j}(g_{k})}}{|C_{G}(g_{k})|}x_{k}$
$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})\overline{\chi_{j}(g_{k})}x_{k}.$
To demonstrate the necessity of Condition 4, notice that if $M_{x}(Q)\geqslant 0$ and $i\in\left[n\right]$, then
$0\leqslant[M_{x}(Q)]_{i1}=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})\overline{\chi_{1}(g_{k})}x_{k}=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})x_{k}.$
Conversely, if Condition 4 holds and $i,j\in[n]$, then
$\displaystyle[M_{x}(Q)]_{ij}$ $\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})\overline{\chi_{j}(g_{k})}x_{k}$
$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})\chi_{\rho_{j}^{*}}(g_{k})x_{k}$ [Eq. 2.3]
$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\rho_{i}\otimes\rho_{j}^{*}}(g_{k})x_{k}$ [Eq. 2.2]
[…]
$\displaystyle=\frac{1}{|G|}\sum_{\ell=1}^{n}\left(\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\langle\chi_{\rho_{i}\otimes\rho_{j}^{*}},\chi_{\ell}\rangle\chi_{\ell}(g_{k})x_{k}\right)$
$\displaystyle=\frac{1}{|G|}\sum_{\ell=1}^{n}\langle\chi_{\rho_{i}\otimes\rho_{j}^{*}},\chi_{\ell}\rangle\left(\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\ell}(g_{k})x_{k}\right)$
[…]
which establishes the sufficiency of Condition 4. ∎
Remark 4.4.
[…]
$\displaystyle M_{x}^{\ast}M_{x}$ $\displaystyle=(QD_{x}D^{-1}Q^{\ast})^{\ast}(QD_{x}D^{-1}Q^{\ast})$
$\displaystyle=(Q\overline{D^{-1}}D_{\overline{x}}Q^{\ast})(QD_{x}D^{-1}Q^{\ast})$
$\displaystyle=Q(\overline{D^{-1}}D_{\overline{x}}D_{x}D^{-1})Q^{\ast}$
$\displaystyle=Q(D_{x}D^{-1}\overline{D^{-1}}D_{\overline{x}})Q^{\ast}$
$\displaystyle=(Q\overline{D^{-1}}D_{\overline{x}}Q^{\ast})(Q\overline{D^{-1}}D_{\overline{x}}Q^{\ast})$
$\displaystyle=M_{x}M_{x}^{\ast}$
[…]
5 Volume of the Projected Spectratope of a Real Character Table
Lemma 5.1.
[…]
scalingConeLemma If $S$ is ideal and $v>0$, then $D_{v}S$ is ideal.
Proof.
Recall that $\mathcal{C}(S)=\mathcal{C}(D_{v}S)$ and if $\alpha_{1},\ldots,\alpha_{n}$ are positive scalars, then $\operatorname{\mathsf{conv}}{(v_{1},\ldots,v_{n})}=\operatorname{\mathsf{conv}}{(\alpha_{1}v_{1},\ldots,\alpha_{n}v_{n})}$. ∎
[…]
If $v_{0},v_{1},\ldots,v_{n}\in\mathbb{R}^{n}$ are affinely independent, i.e., the vectors $v_{1}-v_{0},\ldots,v_{n}-v_{0}$ are linearly independent, then $S:=\operatorname{\mathsf{conv}}(v_{0},v_{1},\ldots,v_{n})$ is called an $n$-simplex. It is known [16] that the volume $V$ of $S$ is given by
[…]
For $k\in\left[n\right]$, let $P_{k}$ the matrix obtained by deleting the $k\textsuperscript{th}$-row of $I_{n}$ and let $\Pi_{k}:\mathbb{F}^{n}\longrightarrow\mathbb{F}^{n-1}$ be the projection map defined by $\Pi_{k}(x)=P_{k}x$.
[…]
Theorem 5.2.
[…]
volumeSpectratope The volume of $\Pi_{1}(\mathcal{P}(D_{v^{-1}}Q))$ is given by
$V=\dfrac{\sqrt{\prod_{k=1}^{n}|C_{G}(g_{k})|}}{(n-1)!\prod_{k=1}^{n}\dim(\rho_{k})}.$ (5.2)
Proof.
Note that
$D_{v^{-1}}=\text{diag}\left(\dfrac{1}{\dim(\rho_{1})},\cdots,\dfrac{1}{\dim(\rho_{n})}\right)$
and
$\det\left(D_{v^{-1}}\right)=\dfrac{1}{\prod_{k=1}^{n}\dim\left(\rho_{k}\right)}.$
Since $Q^{\top}Q$ is a diagonal matrix with $k\textsuperscript{th}$ diagonal entry $|C_{G}(g_{k})|$, it follows that
$\det Q=\sqrt{\prod_{k=1}^{n}|C_{G}(g_{k})|}.$
Applying (5.1) and properties of the determinant,
$\displaystyle V$ $\displaystyle=\dfrac{1}{(n-1)!}\left|\det\left(D_{v^{-1}}Q\right)\right|$
$\displaystyle=\dfrac{1}{(n-1)!}|\det\left(D_{v^{-1}}\right)||\det(Q)|$
$\displaystyle=\dfrac{\sqrt{\prod_{k=1}^{n}|C_{G}(g_{k})|}}{(n-1)!\prod_{k=1}^{n}\dim(\rho_{k})}.\qed$
6 Examples
If $\Lambda=\{\lambda_{1}\coloneqq 1,\lambda_{1},\ldots,\lambda_{n}\}$ is a list of real numbers and $1\leqslant n\leqslant 3$, then $\Lambda$ is realizable if and only if conditions 1.1 and 1.3 are satisfied [13], i.e., if and only if
$\displaystyle 1$ $\displaystyle\geqslant|\lambda_{k}|$ (6.1)
$\displaystyle 1+\sum_{k=1}^{n}\lambda_{k}$ $\displaystyle\geqslant 0.$ (6.2)
Conditions 6.1 and 6.2 define a polytope called the trace nonnegative polytope, whose volume is given by
$2^{n}\left[1-\frac{1}{n!}\sum_{k=0}^{\lfloor\frac{n-1}{2}\rfloor}(-1)^{k}{n\choose k}\left(\frac{n-1}{2}-k\right)^{n}\right].$
(Taylor and Paparella [17, Corollary 2.2]).
Example 6.1.
The character table for $\mathbb{Z}_{2}$ is the Walsh matrix
$H_{2}=\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}$
and the cone $\mathcal{C}(H_{2})$ yields all possible spectra since
$H_{2}D_{x}H_{2}^{-1}=\frac{1}{2}H_{2}D_{x}H_{2}^{\top}=\frac{1}{2}\begin{bmatrix}x_{1}+x_{2}&x_{1}-x_{2}\\ x_{1}-x_{2}&x_{1}+x_{2}\end{bmatrix}$.
Example 6.2.
The character table for $\mathsf{Sym}(3)$ is
$Q=\begin{bmatrix}1&1&1\\ 1&-1&1\\ 2&0&-1\end{bmatrix}$
and
$M_{x}(Q)=\frac{1}{6}\begin{bmatrix}x_{1}+3x_{2}+2x_{3}&x_{1}-3x_{2}+2x_{3}&2(x_{1}-x_{3})\\ x_{1}-3x_{2}+2x_{3}&x_{1}+3x_{2}+2x_{3}&2(x_{1}-x_{3})\\ 2(x_{1}-x_{3})&2(x_{1}-x_{3})&2(2x_{1}+x_{3})\end{bmatrix}$.
[…]
$\displaystyle[M_{x}(Q)]_{33}$
$\displaystyle=$ $\displaystyle\frac{1}{6}\sum_{\ell=1}^{3}\langle\chi_{\rho_{3}\otimes\rho_{3}^{*}},\chi_{\ell}\rangle\left(\sum_{k=1}^{3}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\ell}(g_{k})x_{k}\right)$
$\displaystyle=$ $\displaystyle\frac{1}{6}\left(\frac{4+2}{6}(x_{1}+3x_{2}+2x_{3})+\frac{4+2}{6}(x_{1}-3x_{2}+2x_{3})+\frac{8-2}{6}(2(x_{1}-x_{3}))\right)$
$\displaystyle=$ $\displaystyle\frac{1}{6}(2(2x_{1}+x_{3}))$
which is nonnegative whenever the other three inequalities are nonnegative.
If $v=\begin{bmatrix}1&1&2\end{bmatrix}^{\top}$, then
$D_{v^{-1}}Q=\begin{bmatrix}1&1&1\\ 1&-1&1\\ 1&0&-0.5\end{bmatrix}$
Applying formula 5.2, the area of $\Pi_{1}(\mathcal{P}(D_{v^{-1}}Q))$ is given by
$\displaystyle V$ $\displaystyle=\frac{\sqrt{6\cdot 3\cdot 2}}{(3-1)!\cdot 1\cdot 1\cdot 2}=\frac{3}{2}.$
The area of the trace nonnegative polytope is ${7}/{2}$ so that $\Pi_{1}(\mathcal{P}(D_{v^{-1}}Q))$ occupies ${3}/{7}$ of the feasible region (see Figure 1).
Figure 1: Feasible Region and $\Pi_{1}(\mathcal{P}(D_{v^{-1}}Q))$.
Example 6.3.
The character table of $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ is the Walsh matrix
$H_{4}=\begin{bmatrix}1&1&1&1\\ 1&-1&1&-1\\ 1&1&-1&-1\\ 1&-1&-1&1.\end{bmatrix}$
[…]
$\displaystyle{x_{1}+x_{2}+x_{3}+x_{4}}\geqslant 0$
$\displaystyle{x_{1}-x_{2}+x_{3}-x_{4}}\geqslant 0$
$\displaystyle{x_{1}+x_{2}-x_{3}-x_{4}}\geqslant 0$
$\displaystyle{x_{1}-x_{2}-x_{3}+x_{4}}\geqslant 0$
As $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ is Abelian, $|C_{\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}}(g)|=4,\ \forall g\in\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}.$ Consequently, applying formula 5.2 yields
$\displaystyle V$ $\displaystyle=\frac{\sqrt{4^{4}}}{(3-1)!\cdot 1^{4}}=\frac{8}{3}.$
The volume of the trace nonnegative polytope is ${20}/{3}$ so that $\Pi_{1}(\mathcal{P}(H_{4}))$ occupies ${2}/{5}$ of the feasible region (see Figure 2).
Figure 2: Feasible Region and $\Pi_{1}(\mathcal{P}(H_{4}))$.
Although this polytope only occupies less than half of the realizable spectra, it is known that all possible spectra can be realized via the similarities $H_{4}$ and $H_{2}\oplus H_{2}$ [9, pp. 290–291].
7 Concluding Remarks & Further Inquiry
In what follows, it is assumed, without loss of generality, that $S$ is an ideal Perron similarity normalized such that $Se_{1}=e$ and $|{s_{ij}}|\leqslant 1$ [§§4.3][11]. In such a case, it is known that $\mathcal{C}(S)=\mathcal{C}_{r}(S)$ if and only if $\mathcal{P}(S)=\mathcal{P}_{r}(S)$, where $\mathcal{P}_{r}(S)$ denotes the convex hull of the rows of $S$ [11, Theorem 4.22].
For $x\in\mathbb{C}^{n}$, denote by $\Lambda(x)$ the list $\{x_{1},\dots,x_{n}\}$ and for every natural number $n$, let
$\mathbb{L}^{n}\coloneqq\{x\in\mathbb{C}^{n}\mid\Lambda(x)=\operatorname{\mathsf{spec}}(A),~A\in\mathsf{M}_{n}\left(\mathbb{R}\right),~A\geqslant 0\}$
and
$\mathbb{SL}^{n}\coloneqq\{x\in\mathbb{C}^{n}\mid\Lambda(x)=\operatorname{\mathsf{spec}}(A),~A\in\mathsf{M}_{n}\left(\mathbb{R}\right),~A~\text{stochastic}\}.$
The set $\mathbb{L}^{n}$ is a cone that contains $\mathbb{SL}^{n}$ and a characterization of either set constitutes a solution to the NIEP. It is known that $\mathbb{SL}^{n}$ is compact and star-shaped at $e$ [11, Theorems 3.9 and 3.10].
If $x\in\mathbb{SL}^{n}$, then $1=\rho\left(\Lambda(x)\right)\in\Lambda(x)$ and if $P$ is a permutation matrix, then $\Lambda(Px)=\Lambda(x)$. In light of these two facts, there is no loss in generality in assuming that $x_{1}=1$. It is known that $\Pi_{1}(\mathbb{SL}^{n})$ is star-shaped at the origin [11, Theorem 3.12].
If $x=\begin{bmatrix}1&x_{2}&\cdots&x_{n}\end{bmatrix}^{\top}\in\mathbb{SL}^{n}$, then $x$ is called extremal (in $\mathbb{SL}^{n}$) if $\alpha\Pi_{1}(x)\not\in\Pi_{1}(\mathbb{SL}^{n})$, $\forall\alpha>1$. The set of extremal points in $\mathbb{SL}^{n}$ is denoted by $\mathbb{E}_{n}$. Johnson and Paparella showed that $\mathbb{E}_{n}\subseteq\partial\mathbb{SL}^{n}$ and conjectured the reverse containment. Thus, identifying extremal points is of great interest in the NIEP.
A Perron similarity $S$ is called extremal if $\mathcal{P}(S)$ contains an extremal point other than $e$.
If $G$ is a finite Abelian group, then every irreducible character is one-dimensional and the number of irreducible characters is equal to the order of the group [1, Theorem 10.5.2(a)]. If $\chi_{\rho}$ is a one-dimensional character and $g$ is an element of $G$, then $\chi_{\rho}(g)$ is a power of the primitive root of unity $\omega_{|{g}|}$ [1, p. 303]. Furthermore, since there are prime numbers $p_{1},\dots,p_{k}$ (not necessarily distinct) and positive integers $n_{1},\ldots,n_{k}$ such that
$G\cong\mathbb{Z}_{p_{1}^{n_{1}}}\oplus\cdots\oplus\mathbb{Z}_{p_{k}^{n_{k}}}$
and the discrete Fourier transform matrix $F_{n}$ is the character table of $\mathbb{Z}_{n}$, it follows that
$F_{p_{1}}^{\otimes n_{1}}\otimes\cdots\otimes F_{p_{1}}^{\otimes n_{1}}$
is the character table of $G$. If $G$ has even order, then the character table is of the form
$H_{2^{n_{0}}}\otimes F_{p_{1}}^{\otimes n_{1}}\otimes\cdots\otimes F_{p_{1}}^{\otimes n_{1}}.$
Karpelevič [12] (and previously, Romanovsky [14]) proved that the so-called Karpelevič region
$\Theta_{n}\coloneqq\{\lambda\in\mathbb{C}\mid\lambda\in\operatorname{\mathsf{spec}}A,\ A\geqslant 0,Ae=e\}$
intersects the unit-circle $\{z\in\mathbb{C}\mid|z|=1\}$ at the points $\{\omega_{q}^{p}\mid p/q\in\mathcal{F}_{n}\}$, where $\mathcal{F}_{n}\coloneqq\{p/q\mid 0\leqslant p\leqslant q\leqslant n,~\gcd(p,q)=1\}$ denotes the set of Farey fractions.
Thus, the character table of a finite Abelian group is totally extremal because every entry is extremal in $\Theta_{n}$ (see Figure 2; see Johnson and Paparella [§§8.1][11] for geometrical representations in the four-dimensional complex case).
In view of the above, we offer the following for further inquiry.
Conjecture 7.1.
If $S$ is a normalized ideal Perron similarity and $S$ is totally extremal, then $S$ is the character table of a finite Abelian group.

### v2

@@
 character , character table , group , nonnegative matrix , Perron similarity , polyhedral cone , polytope , representation
-MSC:
-[2020] 15A29 , 20C15 , 20C99 , 15B51 , 52B99
+2020 MSC:
+15A29 , 20C15 , 20C99 , 15B51 , 52B99
+journal: J. Algebra
 […]
@@
 $\mathcal{C}(S):=\{x\in\mathbb{F}^{n}\mid M_{x}\coloneqq SD_{x}S^{-1}\geqslant 0\}\supseteq\{\alpha e\mid\alpha\geqslant 0\},$
-is called the (Perron) spectracone of $S$, and the polytope
+where $D_{x}$ is the diagonal matrix whose $(k,k)$-entry is $x_{k}$, is called the (Perron) spectracone of $S$, and the polytope
 $\mathcal{P}(S):=\left\{x\in\mathcal{C}(S)\mid M_{x}e=e\right\}\supseteq\{e\},$
@@
   • when $m=n$, $\operatorname{\mathsf{spec}}A=\operatorname{\mathsf{spec}}(A)$ denotes the spectrum of $A$ and $\rho=\rho(A)$ denotes the spectral radius of $A$.
-If $X\subseteq\mathbb{F}^{n}$, then the conical hull of $X$, denoted by $\operatorname{\mathsf{coni}}X=\operatorname{\mathsf{coni}}(X)$, is defined by
-$\operatorname{\mathsf{coni}}{X}=\begin{cases}\{0\},&X=\varnothing\\ \left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,~\alpha_{k}\geqslant 0\right\},&X\neq\varnothing.\end{cases}$
-i.e., when $X$ is nonempty, $\operatorname{\mathsf{coni}}X$ consists of all conical combinations.
+If $\varnothing\subset X\subseteq\mathbb{F}^{n}$, then the conical hull of $X$, denoted by $\operatorname{\mathsf{coni}}X=\operatorname{\mathsf{coni}}(X)$, is defined by
+$\operatorname{\mathsf{coni}}{X}=\left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,~\alpha_{k}\geqslant 0\right\}$
+i.e., $\operatorname{\mathsf{coni}}X$ consists of all conical combinations.
 Similarly, the convex hull of $X$, denoted by $\operatorname{\mathsf{conv}}{X}=\operatorname{\mathsf{conv}}(X)$ is defined by
-$\operatorname{\mathsf{conv}}{X}=\begin{cases}\{0\},&X=\varnothing\\ \left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,\ \sum_{k=1}^{m}\alpha_{k}=1,\ \alpha_{k}\geqslant 0\right\},&X\neq\varnothing\end{cases}.$
+$\operatorname{\mathsf{conv}}{X}=\left\{\sum_{k=1}^{m}\alpha_{k}x_{k}\in\mathbb{F}^{n}\mid m\in\mathbb{N},~x_{k}\in X,\ \sum_{k=1}^{m}\alpha_{k}=1,\ \alpha_{k}\geqslant 0\right\}.$
 The conical hull or convex hull of a finite list $\{x_{1},\ldots,x_{n}\}$ is abbreviated to $\operatorname{\mathsf{coni}}(x_{1},\ldots,x_{n})$ or $\operatorname{\mathsf{conv}}(x_{1},\ldots,x_{n})$, respectively.
@@
 Theorem 2.1.
-[…]
-charlinearcombo If $\rho_{1},\ldots,\rho_{n}$ are the distinct irreducible representations of $G$ and $\chi_{1},\ldots,\chi_{n}$ are their respective characters, then
+If $\rho_{1},\ldots,\rho_{n}$ are the distinct irreducible representations of $G$ and $\chi_{1},\ldots,\chi_{n}$ are their respective characters, then
 $\langle\chi_{i},\chi_{j}\rangle=\delta_{ij}.$
@@
 If $x\in\mathbb{R}^{n}$ ($A\in\mathsf{M}_{n}(\mathbb{R})$), then $x$ (respectively, $A$) is viewed as a complex vector (respectively, matrix) via the map $x\longmapsto x+0\mathsf{i}$ (respectively, $A\longmapsto A+0\mathsf{i}$).
+[…]
 If $A\geqslant 0$ and
@@
 Definition 3.1.
-[…]
-spectratope If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then:
+If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then:
   1. (i)
@@
 The set $\mathcal{A}(S)$ is a cone that is closed with respect to matrix multiplication [11, Theorem 4.4(iii)].
-Reall from the introduction that an invertible matrix $S$ is called ideal if $\mathcal{C}(S)=\mathcal{C}_{r}(S)$, where $\mathcal{C}_{r}(S)$ denotes the conical hull of the rows of $S$.
+Recall from the introduction that an invertible matrix $S$ is called ideal if $\mathcal{C}(S)=\mathcal{C}_{r}(S)$, where $\mathcal{C}_{r}(S)$ denotes the conical hull of the rows of $S$.
 It is known that $S$ is ideal if and only if $e\in\mathcal{C}_{r}(S)$ and $r_{i}\in\mathcal{C}(S),\ \forall i\in\left[n\right]$, i.e., every row, viewed as a column vector, belongs to its spectracone [11, Theorem 4.21].
@@
 Definition 3.2.
-[…]
-rhc If $S\in\mathsf{M}_{m\times n}({\mathbb{C}})$, then $S$ is called row Hadamard conic (RHC) if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall i,j\in\left[m\right]$.
+If $S\in\mathsf{M}_{m\times n}({\mathbb{C}})$, then $S$ is called row Hadamard conic (RHC) if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall i,j\in\left[m\right]$.
 Proposition 3.3 ([11, Proposition 4.16]).
-[…]
-simple If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $x^{\top}S^{-1}\geqslant 0$ if and only if $x\in\mathcal{C}_{r}(S)$.
+If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $x^{\top}S^{-1}\geqslant 0$ if and only if $x\in\mathcal{C}_{r}(S)$.
 Proposition 3.4.
-[…]
-rhcrows If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $r_{i}\in\mathcal{C}(S)$ if and only if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),\ \forall j\in\left[n\right]$.
+If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $r_{i}\in\mathcal{C}(S)$ if and only if $r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),\ \forall j\in\left[n\right]$.
 Proof.
@@
 $\displaystyle\Longleftrightarrow\left(r_{j}\circ r_{i}\right)S^{-1}\geqslant 0,~\forall j\in\left[n\right]$
-$\displaystyle\Longleftrightarrow r_{j}\circ r_{i}=r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall j\in\left[n$
+$\displaystyle\Longleftrightarrow r_{j}\circ r_{i}=r_{i}\circ r_{j}\in\mathcal{C}_{r}(S),~\forall j\in\left[n\right]$ [?THM? LABEL:simple]
 as desired. ∎
-[…]
+The following characterization is an immediate result of ?THM? LABEL:rhcrows.
 Theorem 3.5.
-[…]
-RHC ideal If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $S$ is ideal if and only if $e\in\mathcal{C}_{r}(S)$ and $S$ is RHC.
+If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $S$ is ideal if and only if $e\in\mathcal{C}_{r}(S)$ and $S$ is RHC.
 If $S\in\mathsf{GL}_{n}(\mathbb{C})$, then $S$ is called a Perron similarity if there is a diagonal matrix $D$ such that $A=SDS^{-1}$ is irreducible and non-negative.
@@
 Lemma 3.6.
-[…]
-epairconstruct Supoose that $S\in\mathsf{GL}_{n}(\mathbb{C})$, $\lambda\in\mathbb{C}^{n}$, and $A\coloneqq SD_{\lambda}S^{-1}$.
+Supoose that $S\in\mathsf{GL}_{n}(\mathbb{C})$, $\lambda\in\mathbb{C}^{n}$, and $A\coloneqq SD_{\lambda}S^{-1}$.
 If $\exists k\in[n]$ such that $v\coloneqq Se_{k}$ is totally nonzero, then $(\lambda_{k},e)$ is an eigenpair of $D_{v^{-1}}AD_{v}$.
@@
 Theorem 3.7.
-[…]
-P(S)contain If $S\in\mathsf{GL}_{n}(\mathbb{C})$ is a Perron similarity, then there exists $k\in[n]$ such that $\mathcal{P}(S)\subseteq\mathcal{P}(D_{v^{-1}}S)$, where $v\coloneqq Se_{k}$.
+If $S\in\mathsf{GL}_{n}(\mathbb{C})$ is a Perron similarity, then there exists $k\in[n]$ such that $\mathcal{P}(S)\subseteq\mathcal{P}(D_{v^{-1}}S)$, where $v\coloneqq Se_{k}$.
 Proof.
 As $S$ is a Perron similarity, there is a positive integer $k\in[n]$, a nonzero complex number $\alpha$, and a positive vector $x$ such that $v\coloneqq Se_{k}=\alpha x$ is totally nonzero.
-[…]
+If $z\in\mathcal{P}(S)$, then $M_{z}(S)$ is a stochastic matrix.
+Since $(z_{k},v)$ is an eigenpair, it follows that $(z_{k},x)$ is an eigenpair.
+As $M_{z}(S)$ is a nonnegative matrix having a positive eigenvector, it is necessarily the case that $z_{k}=\rho(M_{z}(S))$ [6, Corollary 8.1.30].
+As $M_{z}(S)$ is stochastic, $z_{k}=1$.
+Finally, ?THM?
+LABEL:epairconstruct ensures that $(z_{k},e)=(1,e)$ is an eigenpair of $M_{z}(D_{v^{-1}}S)$.
+∎
 Example 3.8.
-[…]
+The inclusion in ?THM? LABEL:P(S)contain can be strict: if $S=\begin{bmatrix}1&1\\ 2&-2\end{bmatrix}$, then the matrix
 $SD_{x}S^{-1}=\begin{bmatrix}\frac{x_{1}+x_{2}}{2}&\frac{x_{1}-x_{2}}{4}\\ x_{1}-x_{2}&\frac{x_{1}+x_{2}}{2}\end{bmatrix}$
@@
 $\displaystyle=\sum_{k=1}^{n}\left(q_{ik}\left[Q^{-1}\right]_{kj}\right)[e_{1}]_{k}$
-$\displaystyle=\frac{q_{i,1}\overline{q_{j,1}}}{|C_{G}(g_{1})|}$
-$\displaystyle=\frac{\dim(\rho_{i})\dim(\rho_{j})}{|C_{G}(g_{1})|}>0.$
+$\displaystyle=\frac{q_{i,1}\overline{q_{j,1}}}{|C_{G}(e_{G})|}$
+$\displaystyle=\frac{\dim(\rho_{i})\dim(\rho_{j})}{|G|}>0.$
 Thus, $M_{e_{1}}(Q)>0$. ∎
 Theorem 4.2.
+The matrix $Q$ is ideal.
+Proof.
 […]
-extremerayQ The matrix $Q$ is ideal.
-Proof.
-Since
-[…]
-it follows that $e_{1}^{\top}Q=e^{\top}$, i.e., $e\in\mathcal{C}_{r}(Q)$.
-Suppose that $i,j\in[n]$. For ease of notation, let
-$\alpha_{k}\coloneqq\langle\chi_{\rho_{i}\otimes\rho_{j}},\chi_{k}\rangle,\ \forall k\in\left[n\right]$
-and $\alpha\coloneqq\begin{bmatrix}\alpha_{1}&\cdots&\alpha_{n}\end{bmatrix}^{\top}$. With the above in mind,
-$\displaystyle r_{i}^{\top}\circ r_{j}^{\top}$ $\displaystyle=\begin{bmatrix}\chi_{i}(g_{1})\chi_{j}(g_{1})&\cdots&\chi_{i}(g_{n})\chi_{j}(g_{n})\end{bmatrix}$
-$\displaystyle=\begin{bmatrix}\chi_{\rho_{i}\otimes\rho_{j}}(g_{1})&\cdots&\chi_{\rho_{i}\otimes\rho_{j}}(g_{n})\end{bmatrix}$ [Eq. 2.2]
-$\displaystyle=\begin{bmatrix}\sum_{k=1}^{n}\alpha_{k}\chi_{k}(g_{1})&\cdots&\sum_{k=1}^{n}\alpha_{k}\chi_{k}(g_{n})\end{bmatrix}$ [Eq. 2.5]
-$\displaystyle=\alpha^{\top}Q.$
-[…]
-[…]
+The result is clear because the $k$th entry of $r_{i}^{\top}\circ r_{j}^{\top}$ is
+$\chi_{i}(g_{k})\cdot\chi_{j}(g_{k})=\chi_{\rho_{i}\otimes\rho_{j}}(g_{k})=\sum_{\ell=1}^{n}\langle\chi_{\rho_{i}\otimes\rho_{j}},\chi_{\ell}\rangle\chi_{\ell}(g_{k}),$
+which demonstrates that $Q$ is RHC and thus ideal by ?THM? LABEL:RHC_ideal. ∎
+Whereas ?THM?
+LABEL:extremerayQ yields a description of $\mathcal{C}(Q)$ as the conical hull of a set of linearly independent vectors, the subsequent result describes $\mathcal{C}(Q)$ in terms of linear inequalities and demonstrates that not all of the $n^{2}$ inequalities of the matrix equation $M_{x}(Q)\geqslant 0$ are required to specify the cone.
+The linear inequalities are novel sufficient conditions for realizability.
 Theorem 4.3.
-[…]
-ineqredux If $x\in\mathbb{C}^{n}$, then $M_{x}(Q)\geqslant 0$ if and only if
+If $x\in\mathbb{C}^{n}$, then $M_{x}(Q)\geqslant 0$ if and only if
 $\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})x_{k}\geqslant 0,\hskip 10.0pt\forall i\in[n].$ (4.2)
@@
 $\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})\overline{\chi_{j}(g_{k})}x_{k}.$
-To demonstrate the necessity of Condition 4, notice that if $M_{x}(Q)\geqslant 0$ and $i\in\left[n\right]$, then
+To demonstrate the necessity of Condition 4.2, notice that if $M_{x}(Q)\geqslant 0$ and $i\in\left[n\right]$, then
 $0\leqslant[M_{x}(Q)]_{i1}=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})\overline{\chi_{1}(g_{k})}x_{k}=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})x_{k}.$
-Conversely, if Condition 4 holds and $i,j\in[n]$, then
+Conversely, if Condition 4.2 holds and $i,j\in[n]$, then following properties of representations and ?THM? LABEL:charlinearcombo
 $\displaystyle[M_{x}(Q)]_{ij}$ $\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})\overline{\chi_{j}(g_{k})}x_{k}$
-$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{i}(g_{k})\chi_{\rho_{j}^{*}}(g_{k})x_{k}$ [Eq. 2.3]
-$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\rho_{i}\otimes\rho_{j}^{*}}(g_{k})x_{k}$ [Eq. 2.2]
-[…]
+$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\rho_{i}\otimes\rho_{j}^{*}}(g_{k})x_{k}$
+$\displaystyle=\frac{1}{|G|}\sum_{k=1}^{n}\left(\sum_{\ell=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\langle\chi_{\rho_{i}\otimes\rho_{j}^{*}},\chi_{\ell}\rangle\chi_{\ell}(g_{k})x_{k}\right)$
 $\displaystyle=\frac{1}{|G|}\sum_{\ell=1}^{n}\left(\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\langle\chi_{\rho_{i}\otimes\rho_{j}^{*}},\chi_{\ell}\rangle\chi_{\ell}(g_{k})x_{k}\right)$
-$\displaystyle=\frac{1}{|G|}\sum_{\ell=1}^{n}\langle\chi_{\rho_{i}\otimes\rho_{j}^{*}},\chi_{\ell}\rangle\left(\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\ell}(g_{k})x_{k}\right)$
-[…]
-which establishes the sufficiency of Condition 4. ∎
+$\displaystyle=\frac{1}{|G|}\sum_{\ell=1}^{n}\langle\chi_{\rho_{i}\otimes\rho_{j}^{*}},\chi_{\ell}\rangle\left(\sum_{k=1}^{n}|\operatorname{\mathsf{cl}}(g_{k})|\chi_{\ell}(g_{k})x_{k}\right)\geqslant 0,$
+which establishes the sufficiency of Condition 4.2. ∎
 Remark 4.4.
@@
 $\displaystyle=(Q\overline{D^{-1}}D_{\overline{x}}Q^{\ast})(QD_{x}D^{-1}Q^{\ast})$
-$\displaystyle=Q(\overline{D^{-1}}D_{\overline{x}}D_{x}D^{-1})Q^{\ast}$
-$\displaystyle=Q(D_{x}D^{-1}\overline{D^{-1}}D_{\overline{x}})Q^{\ast}$
+$\displaystyle=Q(\overline{D^{-1}}D_{\overline{x}}DD_{x}D^{-1})Q^{\ast}$
+$\displaystyle=Q(D_{x}D^{-1}D\overline{D^{-1}}D_{\overline{x}})Q^{\ast}$
 $\displaystyle=(Q\overline{D^{-1}}D_{\overline{x}}Q^{\ast})(Q\overline{D^{-1}}D_{\overline{x}}Q^{\ast})$
@@
 Lemma 5.1.
-[…]
-scalingConeLemma If $S$ is ideal and $v>0$, then $D_{v}S$ is ideal.
+If $S$ is ideal and $v>0$, then $D_{v}S$ is ideal.
 Proof.
@@
 ∎
-[…]
+If $v\coloneqq Qe_{1}$, then, since $Q$ is ideal, it follows that $D_{v^{-1}}Q$ is ideal and $(D_{v^{-1}}Q)e_{1}=e$.
+Recall, by ?THM?
+LABEL:P(S)contain, that $\mathcal{P}(Q)\subseteq\mathcal{P}(D_{v^{-1}}Q)$.
 If $v_{0},v_{1},\ldots,v_{n}\in\mathbb{R}^{n}$ are affinely independent, i.e., the vectors $v_{1}-v_{0},\ldots,v_{n}-v_{0}$ are linearly independent, then $S:=\operatorname{\mathsf{conv}}(v_{0},v_{1},\ldots,v_{n})$ is called an $n$-simplex.
@@
 Theorem 5.2.
-[…]
-volumeSpectratope The volume of $\Pi_{1}(\mathcal{P}(D_{v^{-1}}Q))$ is given by
+The volume of $\Pi_{1}(\mathcal{P}(D_{v^{-1}}Q))$ is given by
 $V=\dfrac{\sqrt{\prod_{k=1}^{n}|C_{G}(g_{k})|}}{(n-1)!\prod_{k=1}^{n}\dim(\rho_{k})}.$ (5.2)
@@
 $M_{x}(Q)=\frac{1}{6}\begin{bmatrix}x_{1}+3x_{2}+2x_{3}&x_{1}-3x_{2}+2x_{3}&2(x_{1}-x_{3})\\ x_{1}-3x_{2}+2x_{3}&x_{1}+3x_{2}+2x_{3}&2(x_{1}-x_{3})\\ 2(x_{1}-x_{3})&2(x_{1}-x_{3})&2(2x_{1}+x_{3})\end{bmatrix}$.
-[…]
+By ?THM?
+LABEL:ineqredux, the half-spaces that determine the polyhedral cone $\mathcal{C}(Q)$ are $x_{1}+3x_{2}+2x_{3}\geqslant 0$, $x_{1}-3x_{2}+2x_{3}\geqslant 0$, and $2(x_{1}-x_{3})\geqslant 0$.
+Indeed, the inequality $2(2x_{1}+x_{3})\geqslant 0$ is redundant since
 $\displaystyle[M_{x}(Q)]_{33}$
@@
 $H_{4}=\begin{bmatrix}1&1&1&1\\ 1&-1&1&-1\\ 1&1&-1&-1\\ 1&-1&-1&1.\end{bmatrix}$
-[…]
+If $x\in\mathbb{R}^{4}$, then, following ?THM? LABEL:ineqredux, $M_{x}(H_{4})\geqslant 0$ if and only if
 $\displaystyle{x_{1}+x_{2}+x_{3}+x_{4}}\geqslant 0$
@@
 As $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ is Abelian, $|C_{\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}}(g)|=4,\ \forall g\in\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}.$ Consequently, applying formula 5.2 yields
-$\displaystyle V$ $\displaystyle=\frac{\sqrt{4^{4}}}{(3-1)!\cdot 1^{4}}=\frac{8}{3}.$
+$\displaystyle V$ $\displaystyle=\frac{\sqrt{4^{4}}}{(4-1)!\cdot 1^{4}}=\frac{8}{3}.$
 The volume of the trace nonnegative polytope is ${20}/{3}$ so that $\Pi_{1}(\mathcal{P}(H_{4}))$ occupies ${2}/{5}$ of the feasible region (see Figure 2).